If barium has a threshold frequency of 6.07 x 1014 Hz, calculate the kinetic energy of the electron emitted when barium is illuminated with a wavelength of 385 nm. 


1
Expert's answer
2021-04-20T17:30:18-0400

"E=h(\frac{c}{\lambda}-\nu),"

"E=6.63\cdot 10^{-34}\cdot(\frac{3\cdot 10^8}{385\cdot10^{-9}}-6.07\cdot 10^{-14})=11.4\cdot 10^{-20}~J."